Minutes of the BEAST meeting
Friday July 10th 12:00 JST

  • Structure:
    The drift tube quadrants are mechanically complete and assembly will be finished shortly (2/3 complete). The structure was shown at the meeting (see photo on the BEAST WEB page). The design of the SVD quadrant is 80% finished. The drift tubes for the outer BEAST are being fabricated. The design for the beampipe support and the BEAST support still needs to be finalized. A first prototype was shown at the meeting.

    A shelf for the CsI(Tl) crystals was manufactured and shown at the meeting. It supports the crystal and allows the angle of inclination of the crystal to be adjusted. The radii of the outer endcap CsI(Tl) crystals need to be fixed. The question of whether thick Pb shielding for the CsI(Tl) crystals is needed was discussed. These items will be decided soon by Boris Shwartz and the CsI group.

    The 32 calibrated MOSFETS will be shipped from Australia to Hawaii. Cool dry air from an air conditioner to keep the temperature of the diodes constant will be needed at KEK. A simple gas system with distributors and bubblers as well as a supply of HRS gas (Ar 89%, Co2 10% CH4 1%) will also needed at KEK.

  • Drift tubes:
    Now checking the tension of the drift tube wires (by using magnetic induction and an acoustic method). 10 of the PC Readout boards are completely populated and the rest should be finished soon. Recall there are AC and DC coupled outputs from the drift tube PC board.
    Sahu et al have checked that the DC performance is linear. Gary Varner found a bug in the design for the AC part - a jumper has been added and this part now works.

  • Diodes:
    The temperature calibration has been done for 8 diodes in the range 20-35 C. As expected, the temperature dependence vs 1/T is exponential so we will need air conditioning to keep the calibration stable. The depletion depth looks ok. The diodes will be operated in the range 50-60 V. The PC readout for the diodes also has an AC and DC coupled outputs. The DC part works. The AC part has a bug which should be corrected in the final production run. The radiation dose calibration for the new batch of diodes will begin shortly.

  • DAQ:
    3 CAMAC modules were carried to Australia last week. A Kinetics 3922 crate controller will be sent to Melbourne from KEK. The interface board to the PC is the remaining item to be decided (either a Kinetics 2927 or 2915 board) and will have to be purchased.

  • Simulation:
    Keith Hamasaki showed the status of the GEANT simulation which is now done using the standard BELLE gsim framework. The CsI(Tl) crystals and drift tubes have been added. Trial runs with the background generator were tried. The background generator, however, still uses a lattice from last year.